OpenStudy (rositaluv):

a^4 / a^2 What would this be in standard form?

jimthompson5910 (jim_thompson5910):

The bases are the same, so subtract the exponents.

jimthompson5910 (jim_thompson5910):

Answer is a^k where k is the result from that subtraction

OpenStudy (rositaluv):

a^2

jimthompson5910 (jim_thompson5910):

yes, that's the correct answer

OpenStudy (rositaluv):

what would it be in expanded form?

OpenStudy (rositaluv):

That is the question i meant to post

jimthompson5910 (jim_thompson5910):

a^4 = a*a*a*a a^2 = a*a

OpenStudy (rositaluv):

Thanks so much

jimthompson5910 (jim_thompson5910):

You can then see that 2 'a' terms cancel from each